# A Survey of LLM-based Automated Program Repair: Taxonomies, Design Paradigms, and Applications

### v1

@@ 4. Procedural Approaches
Procedural approaches treat APR as a fixed pipeline where the language model runs only at scripted checkpoints and never alters its weights or the surrounding logic. Every choice about search scope, context retrieval, validation order, and iteration budget is hard-coded, which yields reproducible outcomes and predictable cost. Within this umbrella, four recurring subclasses can be identified. Test-in-the-loop pipelines alternate generation with unit tests and take each failure as an oracle (Xia and Zhang, 2024; Yin et al., 2024; Tang et al., 2024; Kong et al., 2024). Human-in-the-loop pipelines thread developer insight into the loop through hints, plans, or reviews that shape the next prompt (Yang et al., 2024c; Takerngsaksiri et al., 2024; Zhao et al., 2024b). Retrieval-augmented pipelines, or RAG-in-the-loop, insert concise evidence fetched by static rules between calls (Xia et al., 2024; Mansur et al., 2024; Yang et al., 2025a). Analysis-augmented pipelines, or AAG-in-the-loop, invoke compilers, debuggers, or security analyzers to feed diagnostics or traces that steer regeneration toward the root cause (Brancas et al., 2025; Xiao et al., 2025; Fakih et al., 2025; Wei et al., 2023). These subclasses differ mainly in the source of the feedback they inject and in the auxiliary tools that collect it. At the same time, they all rely on scripted control to keep the repair deterministic.
@@ 4.1. Test-in-the-Loop Pipelines
Test-in-the-loop pipelines keep an LLM frozen and alternate patch generation with automatic test execution. Lightweight conversational loops, exemplified by ChatRepair (Xia and Zhang, 2024), feed failing assertions back to the LLM and regenerate patches in the loop. Reasoning-enhanced variants such as ThinkRepair (Yin et al., 2024) cache a chain of thought and partial passes. Bandit-guided search, represented by REx (Tang et al., 2024), allocates queries based on test rewards, whereas contrast-aware schemes like ContrastRepair (Kong et al., 2024) pair failing and passing tests to isolate the faulty logic. Because every step and data flow is predetermined by the script rather than the model, these pipelines differ from fully agentic frameworks while still turning the test suite into an iterative oracle that steers the LLM toward a correct fix.
ChatRepair prompts GPT-3.5 or GPT-4 to propose a patch, executes the full test suite, then feeds the exact failing assertion messages plus the buggy lines back to the model for up to five regeneration rounds (Xia and Zhang, 2024). Its conversational prompts are lightweight and cost-efficient, letting ChatGPT fix 162/337 Defects4J bugs at roughly \$0.42 per bug while requiring no auxiliary analysis. Strengths include minimal engineering effort and strong empirical performance on real Java defects. Weaknesses include prompt sensitivity and limited scalability to multi-hunk or cross-file bugs, as failure messages alone may not pinpoint the root cause. All loop conditions, timeouts, and data exchanges are hard-wired; the LLM only supplies patch text and never alters the surrounding script.
@@
ContrastRepair augments conversation-based APR by supplying ChatGPT with a contrastive pair of nearly identical test cases (one failing and one passing) to focus the model on the specific code change that triggers the bug (Kong et al., 2024). It repairs 143 of 337 bugs on Defects4J (Just et al., 2014), and it also achieves the best results on QuixBugs (Lin et al., 2017) and HumanEval-Java (Jiang et al., 2023a). The authors demonstrate that selecting the passing test with the smallest Damerau-Levenshtein distance (Damerau, 1964; Levenshtein et al., 1966) from the failing test yields richer feedback and reduces the number of conversation rounds required for a correct patch. However, this method relies on the presence of at least one closely matching passing test and extra retrieval time, so its benefit diminishes when no closely matching passing case exists, when generating one by mutation exceeds the time budget, or when the synthetic passing test hides the real fault.
These representatives illustrate two consistent patterns. First, richer feedback yields higher fix rates: moving from raw assertion messages to contrastive test pairs, chain-of-thought traces, or adaptive candidate selection progressively boosts success without altering model weights. Second, feedback granularity trades off against resource usage. Lightweight conversations scale to hundreds of bugs at minimal cost, whereas trace-driven, contrast-driven, or search-intensive loops achieve greater accuracy at the expense of longer prompts, more LLM calls, and added engineering complexity. A promising future direction is an adaptive tiered pipeline that escalates from inexpensive cues to detailed analysis only when simpler signals fail, balancing efficiency with effectiveness across diverse defect types.
@@ 4.2. Human-in-the-Loop Pipelines
Earlier non-LLM studies have underscored the necessity of a human in the loop for APR (Geethal et al., 2023; Sobania et al., 2023). Human-in-the-loop pipelines embed frozen LLMs in a structured workflow that alternates between the LLMs’ inference and human insight. Recent work spans interactive tutoring for novice programmers (Yang et al., 2024c), planner-and-coder collaborations on industrial issues (Takerngsaksiri et al., 2024), and reviewer-driven patch refinement (Zhao et al., 2024b). For each human-in-the-loop approach, progress occurs when the LLM revises its patch after receiving new human feedback, such as tutor hints, input labels, design rationales, or reviewer advice.
CREF partners a programming tutor with a frozen GPT-4 in an interactive repair dialogue (Yang et al., 2024c). Each cycle shows the model the buggy function, failing tests, and the tutor’s natural-language hint. The tutor then reviews the candidate’s patch, supplying a fresh hint or approval that becomes part of the next prompt. Because new human guidance is injected at every turn, the dialogue converges quickly and lifts top-5 correctness to 76.6% on the 1,239-bug TutorCode (TutorCode, 2024) benchmark, 17-25% above no-tutor baselines.
HULA is a human-in-the-loop LLM agent framework, where an AI Planner and an AI Coding Agent work with a human engineer to localize files, draft a coding plan, and generate code patches for each issue (Takerngsaksiri et al., 2024). Experimental results on SWE-bench, the planner reached 86% file-recall and the coder attained 45% LLM-based code-similarity, while in an online study with 663 real Jira issues 82% of plans were approved and 56 pull requests were merged. The authors find that detailed issue descriptions and interactive human feedback are critical, since practitioner agreement rises when engineers enrich descriptions and review agent output. Nevertheless, HULA still struggles to guarantee functional correctness beyond unit tests and requires substantial context preparation.
DRCodePilot embeds frozen GPT-4 in a dialogue-style loop that repeatedly injects developer design rationales and reviewer-style feedback (Zhao et al., 2024b). Each cycle starts with DRMiner (Zhao et al., 2024c)-extracted solution information, letting GPT-4 locate the buggy segment and draft a patch; a tuned CodeT5P (Wang et al., 2023) then supplies a reference patch, while an identifier recommender offers renaming hints. GPT-4 reflects on this mixed human feedback, producing a refined fix. On a new 938-issue Flink and Solr benchmark, it yields 109 and 18 exact-match patches, respectively, 4.7x and 3.6x more than GPT-4 zero-shot, with CodeBLEU gains of 5.4% and 3.9%.
@@ 4.3. RAG-in-the-Loop Pipelines
RAG-in-the-Loop pipelines keep the LLM frozen and insert a deterministic retrieval stage between successive generations so that each new prompt is enriched with fresh, highly targeted evidence. The control flow is scripted by the designer rather than chosen by the model, placing these methods squarely in the procedural paradigm. Within this paradigm, hierarchical repository pruning exemplified by Agentless (Xia et al., 2024), forum-augmented iteration as used in RAGFix (Mansur et al., 2024), and knowledge-graph expansion adopted in KGCompass (Yang et al., 2025a) all show that precise retrieval can raise repair accuracy while containing token cost.
Agentless follows a three-phase loop that first asks GPT-4o to rank suspicious files, then shows only their signatures so the model can pick concrete edit lines, and finally reveals full code for patch generation, followed by test validation (Xia et al., 2024). The hierarchical retrieval supplies exactly the required context at each stage and avoids exposing the entire repository up front. On SWE-bench Lite, the initial pipeline with GPT-4o repaired 32.0% of issues at about \$0.70 per bug, outperforming many tool-driven autonomous agents with much less tokens. More recently, Agentless has been adopted by various leading companies (including OpenAI, Meta, and DeepSeek) and inspired the design of various other recent software agents. While promising, the repository traversal can be inefficient or imprecise on very large projects, and the fixed three rounds may under-explore complex defects, suggesting adaptive depth or early-exit heuristics.
RAGFix targets Python errors by mining a six-million-post Stack Overflow vector knowledge graph: after an initial failure, it extracts the stack trace, retrieves the most relevant discussions, summarizes them, and prepends the advice to a follow-up GPT-3.5 or Llama-3 prompt (Mansur et al., 2024). The system iterates once or twice, each cycle combining feature extraction, retrieval, and regeneration, which markedly raises pass@1 accuracy on the HumanEvalFix (Muennighoff et al., 2023) benchmark, especially for small 8b-parameter models that lack strong built-in API knowledge. Leveraging real developer discussions provides rich, domain-specific hints without any parameter updates. Effectiveness depends on high-quality forum matches; therefore, projects that rely on proprietary APIs may require an in-house corpus or hybrid retrieval sources.
@@ 4.4. AAG-in-the-Loop Pipelines
Analysis-augmented generation (AAG) pipelines interleave an automated program analysis pass with LLM repair steps. Evidence can stem from logical inconsistency sets exploited by FormHe (Brancas et al., 2025), from semantics-based token pruning in Repilot (Wei et al., 2023), or static and hybrid security analyzers such as CodeQL (De Moor et al., 2007) adopted by PredicateFix (Xiao et al., 2025) and the exploit validators embedded in LLM4CVE (Fakih et al., 2025). These signals guide regeneration without relying solely on unit tests.
@@
Repilot leverages a frozen LLM with a semantics-based Completion Engine—concretely, the incremental parser and code-completion API of Eclipse JDT Language Server (Foundation, 2023; Foundation and Wei, 2023), which inspects the partial AST, enumerates type-, scope-, and import-consistent continuations, prunes invalid tokens, and actively completes rare identifiers during autoregressive decoding (Wei et al., 2023). The workflow iterates token by token: the LLM supplies a probability list for the next token, the engine eliminates options statically infeasible, and occasionally inserts a full identifier when only one continuation type-checks, after which generation resumes. Evaluated on single-hunk Java bugs from Defects4J 1.2/2.0, Repilot fixes 66 and 50 bugs, respectively—27 % and 47 % more than prior state-of-the-art APR tools—and sustains a 66 % compilation rate at 30 samples per bug. Because static analysis is consulted for every partial program, Repilot delivers markedly higher validity and unique fixes but incurs roughly 7 % overhead for CodeT5-large (Wang et al., 2021) and negligible overhead for InCoder-6.7B (Fried et al., 2022). The authors argue that such fine-grained pruning is orthogonal to model choice and can be generalized to other generation tasks. However, scalability to multi-file projects and languages beyond Java remains an open question.
FormHe combines an ASP solver with a fine-tuned LLM for student assignments (Brancas et al., 2025). The solver first computes Minimal Strongly Inconsistent Correction Subsets and other hints, identifying the rules that violate the reference specification. These hints form a compact prompt for the LLM, which rewrites only the suspicious rules. If the repaired program still produces extra or missing answer sets, the solver extracts new MSICSs (Kullmann, 2009) and the loop repeats. In classroom data, the tool pinpointed faulty lines in $94\%$ of submissions and produced fully correct repairs for $58\%$ of the incorrect cases.
@@
PredicateFix targets security alerts raised by CodeQL (De Moor et al., 2007) and GoInsight. Their analyzer inspects the rule predicates behind an alert, derives a bridging predicate that links the defect to missing sanitization, and then retrieves key examples that satisfy that predicate (Xiao et al., 2025). These examples, along with the original alert, are incorporated into a retrieval-augmented prompt, guiding GPT-4 or other models to generate a patch. After each proposal, the analyzer reruns; unresolved alerts trigger another round while already-fixed code is frozen. Across 6,027 CVE-based alerts, the pipeline raised the number of correct repairs by $27.1\%$-$72.5\%$ relative to plain RAG baselines.
@@
4.5. Discussion and Outlook
Procedural workflows provide predictable cost and reproducible behavior because the script predetermines every retrieval step, test execution, and human intervention. Results from ChatRepair, ThinkRepair, Agentless, and KGCompass indicate that repair accuracy rises as the feedback given to a frozen model becomes more precise; structured traces, knowledge-graph context, and reviewer comments offer clearer guidance than raw assertion text. Future work can improve efficiency by prioritizing feedback according to its information value, so that prompts contain only the most helpful facts, and by allowing the loop to increase its analytical depth only when simpler cues are insufficient. Integrating firm control of scripted pipelines with small autonomous routines for complex faults may extend coverage without incurring additional costs. These advances will support the reliable adoption of LLM-based repair in continuous integration practices.

### v2

@@ 6. Procedural Approaches
Procedural approaches treat software repair as a fixed pipeline where the language model runs only at scripted checkpoints and never alters its weights or the surrounding logic. Every choice about search scope, context retrieval, validation order, and iteration budget is hard-coded, which yields reproducible outcomes and predictable costs. Within this umbrella, four recurring subclasses can be identified. Test-in-the-loop pipelines alternate generation with unit tests and take each failure as an oracle (Xia and Zhang, 2024; Yin et al., 2024; Tang et al., 2024; Kong et al., 2025). Human-in-the-loop pipelines thread developer insight into the loop through hints, plans, or reviews that shape the next prompt (Yang et al., 2024b; Takerngsaksiri et al., 2025; Zhao et al., 2024b). Retrieval-augmented pipelines, or RAG-in-the-loop, insert concise evidence fetched by static rules between calls (Xia et al., 2025; Fei et al., 2025; Yang et al., 2025a). Analysis-augmented pipelines, or AAG-in-the-loop, invoke compilers, debuggers, or security analyzers to feed diagnostics or traces that steer regeneration toward the root cause (Kim et al., 2025; Xiao et al., 2025; Fakih et al., 2025; Wei et al., 2023). These subclasses differ mainly in the source of the feedback they inject and in the auxiliary tools that collect it. At the same time, they all rely on scripted control to keep the repair deterministic.
@@ 6.1. Test-in-the-Loop Pipelines
Test-in-the-loop pipelines keep an LLM frozen and alternate patch generation with automatic test execution. Lightweight conversational loops, exemplified by ChatRepair (Xia and Zhang, 2024), feed failing assertions back to the LLM and regenerate patches in the loop. Reasoning-enhanced variants such as ThinkRepair (Yin et al., 2024) cache a chain of thought and partial passes. Bandit-guided search, represented by REx (Tang et al., 2024), allocates queries based on test rewards, whereas contrast-aware schemes, such as ContrastRepair (Kong et al., 2025), pair failing and passing tests to isolate the faulty logic. Because every step and data flow is predetermined by the script rather than the model, these pipelines differ from fully agentic frameworks while still turning the test suite into an iterative oracle that steers the LLM toward a correct fix.
ChatRepair prompts GPT-3.5 or GPT-4 to propose a patch, executes the full test suite, then feeds the exact failing assertion messages plus the buggy lines back to the model for an average of 21.86 regeneration rounds (Xia and Zhang, 2024). Its conversational prompts are lightweight and cost-efficient, letting ChatGPT fix 162/337 Defects4J bugs at roughly \$0.42 per bug while requiring no auxiliary analysis. Strengths include minimal engineering effort and strong empirical performance on real Java defects. Weaknesses include prompt sensitivity and limited scalability to multi-hunk or cross-file bugs, as failure messages alone may not pinpoint the root cause. All loop conditions, timeouts, and data exchanges are hard-wired; the LLM only supplies patch text and never alters the surrounding script.
@@
ContrastRepair augments conversation-based software repair by supplying ChatGPT with a contrastive pair of nearly identical test cases (one failing and one passing) to focus the model on the specific code change that triggers the bug (Kong et al., 2025). It repairs 143 of 337 bugs on Defects4J (Just et al., 2014), and it also achieves the best results on QuixBugs (Lin et al., 2017) and HumanEval-Java (Jiang et al., 2023a). The authors demonstrate that selecting the passing test with the smallest Damerau-Levenshtein distance (Damerau, 1964; Levenshtein et al., 1966) from the failing test yields richer feedback and reduces the number of conversation rounds required for a correct patch. However, this method relies on the presence of at least one closely matching passing test and extra retrieval time, so its benefit diminishes when no closely matching passing case exists, when generating one by mutation exceeds the time budget, or when the synthetic passing test hides the real fault.
These representatives illustrate two consistent patterns. First, richer feedback yields higher fix rates: moving from raw assertion messages to contrastive test pairs, chain-of-thought traces, or adaptive candidate selection progressively boosts success without altering model weights. Second, feedback granularity trades off against resource usage. Lightweight conversations scale to hundreds of bugs at minimal cost. In contrast, trace-driven, contrast-driven, or search-intensive loops achieve greater accuracy at the expense of longer prompts, more LLM calls, and added engineering complexity. A promising future direction is an adaptive tiered pipeline that escalates from inexpensive cues to detailed analysis only when simpler signals fail, balancing efficiency with effectiveness across diverse defect types.
@@ 6.2. Human-in-the-Loop Pipelines
Earlier non-LLM studies have underscored the necessity of a human in the loop for software repair (Geethal et al., 2023; Sobania et al., 2023). Human-in-the-loop pipelines embed frozen LLMs into a structured workflow that alternates between LLM inference and human insight. Recent work spans interactive tutoring for novice programmers (Yang et al., 2024b), planner-and-coder collaborations on industrial issues (Takerngsaksiri et al., 2025), and reviewer-driven patch refinement (Zhao et al., 2024b). For each human-in-the-loop approach, progress occurs when the LLM revises its patch after receiving new human feedback, such as tutor hints, input labels, design rationales, or reviewer advice.
CREF partners a programming tutor with a frozen GPT-4 in an interactive repair dialogue (Yang et al., 2024b). Each cycle shows the model the buggy function, failing tests, and the tutor’s natural-language hint. The tutor then reviews the candidate’s patch, supplying a fresh hint or approval that becomes part of the next prompt. Because new human guidance is injected at every turn, the dialogue converges quickly and lifts top-5 correctness to 76.6% on the 1,239-bug TutorCode (TutorCode, 2024) benchmark, 17-25% above no-tutor baselines.
HULA is a human-in-the-loop LLM agent framework, where an AI Planner and an AI Coding Agent work with a human engineer to localize files, draft a coding plan, and generate code patches for each issue (Takerngsaksiri et al., 2025). Experimental results on SWE-bench, the planner reached 86% file-recall and the coder attained 45% LLM-based code-similarity, while in an online study with 663 real Jira issues 82% of plans were approved and 56 pull requests were merged. The authors find that detailed issue descriptions and interactive human feedback are critical, since practitioner agreement rises when engineers enrich descriptions and review agent output. Nevertheless, HULA still struggles to guarantee functional correctness beyond unit tests and requires substantial context preparation.
DRCodePilot embeds frozen GPT-4 in a dialogue-style loop that repeatedly injects developer design rationales and reviewer-style feedback (Zhao et al., 2024b). Each cycle starts with DRMiner (Zhao et al., 2024c)-extracted solution information, letting GPT-4 locate the buggy segment and draft a patch; a tuned CodeT5P (Wang et al., 2023) then supplies a reference patch, while an identifier recommender offers renaming hints. GPT-4 reflects on this mixed human feedback and produces a refined fix. On a new 938-issue Flink and Solr benchmark, it yields 109 and 18 exact-match patches, respectively, 4.7x and 3.6x more than GPT-4 zero-shot, with CodeBLEU gains of 5.4% and 3.9%.
@@ 6.3. RAG-in-the-Loop Pipelines
RAG-in-the-Loop pipelines keep the LLM frozen and insert a deterministic retrieval stage between successive generations, enriching each new prompt with fresh, highly targeted evidence. The control flow is scripted by the designer rather than chosen by the model, placing these methods squarely in the procedural paradigm. Within this paradigm, hierarchical repository pruning exemplified by Agentless (Xia et al., 2025), multi-stage retrieving as used in PATCH (Zhang et al., 2025), and knowledge-graph expansion adopted in KGCompass (Yang et al., 2025a) all show that precise retrieval can raise repair accuracy while containing token cost.
Agentless follows a three-phase loop that first asks GPT-4o to rank suspicious files, then shows only their signatures so the model can pick concrete edit lines, and finally reveals full code for patch generation, followed by test validation (Xia et al., 2025). Hierarchical retrieval provides exactly the required context at each stage and avoids exposing the entire repository up front. On SWE-bench Lite, the initial pipeline with GPT-4o repaired 32.0% of issues at about \$0.70 per bug, outperforming many tool-driven autonomous agents with much fewer tokens. While promising, the repository traversal can be inefficient or imprecise on very large projects, and the fixed three rounds may under-explore complex defects, suggesting adaptive depth or early-exit heuristics.
PATCH is a stage-wise framework that augments the buggy code snippet with class and repository-level dependence context, commit message programmer intent, and retrieved bug-fixing demonstrations, then simulates tester, developer, and reviewer roles with three ChatGPT agents across bug reporting, bug diagnosis, patch generation, and patch verification (Zhang et al., 2025). On the BFP (Tufano et al., 2019) benchmark, PATCH reaches pass@1 33.97%, pass@3 37.08%, and pass@5 39.81% with a Levenshtein distance of 21.44, compared with GPT-4 at 19.96%, 24.42%, 26.00% and 26.07 respectively. For RAG in the loop program repair, the main lesson is that retrieval of dependence context, commit messages, and similar bug fixing pairs is most effective when combined with an explicit multi-role, multi-stage interaction loop that lets the LLM iteratively explain, reason, and review patches instead of treating repair as a single pass generation task.
@@ 6.4. AAG-in-the-Loop Pipelines
Analysis-augmented generation (AAG) pipelines interleave an automated program analysis pass with LLM repair steps. Evidence can stem from sanitizer logs analyzed by SAN2PATCH (Kim et al., 2025), from semantics-based token pruning in Repilot (Wei et al., 2023), or from static and hybrid security analyzers such as CodeQL (De Moor et al., 2007), adopted by PredicateFix (Xiao et al., 2025), and from exploit validators embedded in LLM4CVE (Fakih et al., 2025). A fixed script orchestrates the analysis stages and regeneration steps, keeping these pipelines within the procedural paradigm. The analysis signals guide regeneration without relying solely on unit tests.
@@
Repilot leverages a frozen LLM with a semantics-based Completion Engine—concretely, the incremental parser and code-completion API of Eclipse JDT Language Server (Foundation, 2023; Foundation and Wei, 2023), which inspects the partial AST, enumerates type-, scope-, and import-consistent continuations, prunes invalid tokens, and actively completes rare identifiers during autoregressive decoding (Wei et al., 2023). The workflow iterates token by token: the LLM supplies a probability list for the next token, the engine eliminates options statically infeasible, and occasionally inserts a full identifier when only one continuation type-checks, after which generation resumes. Evaluated on single-hunk Java bugs from Defects4J 1.2/2.0, Repilot fixes 66 and 50 bugs, respectively, 27% and 47% more than prior state-of-the-art repair tools, and sustains a 59% compilation rate at 5000 samples per bug. Because static analysis is consulted for every partial program, Repilot delivers markedly higher validity and unique fixes but incurs roughly 7% overhead for CodeT5-large (Wang et al., 2021) and negligible overhead for InCoder-6.7B (Fried et al., 2022). The authors argue that such fine-grained pruning is orthogonal to model choice and can be generalized to other generation tasks. However, scalability to multi-file projects and languages beyond Java remains an open question.
SAN2PATCH is an automated vulnerability repair system that uses only sanitizer logs and source code, running a four-stage Tree of Thought LLM pipeline (Vulnerability Comprehension, Where-To-Fix, How-To-Fix, Candidate Generation) with AST-based code context retrieval and strict patch validation to synthesize and verify patches (Kim et al., 2025). On the VulnLoc (Shen et al., 2021) benchmark, SAN2PATCH with gpt-4o correctly repairs 31 of 39 vulnerabilities, achieving a 79.5% success rate. For AAG-in-the-loop vulnerability repair, the key insight is that careful prompt structuring and multi-stage LLM reasoning over lightweight runtime artifacts like sanitizer logs can replace heavy program analysis and manual localization while still producing functionally correct patches at scale.
@@
PredicateFix targets security alerts raised by CodeQL (De Moor et al., 2007) and GoInsight. Their analyzer inspects the rule predicates behind an alert, derives a bridging predicate that links the defect to missing sanitization, and then retrieves key examples that satisfy that predicate (Xiao et al., 2025). These examples, along with the original alert, are incorporated into a retrieval-augmented prompt that guides GPT-4 or other models to generate a patch. After each proposal, the analyzer reruns; unresolved alerts trigger another round while already-fixed code is frozen. Across 6,027 CVE-based alerts, the pipeline raised the number of correct repairs by $27.1\%$-$72.5\%$ relative to plain RAG baselines.
@@
Overview Procedural workflows provide predictable cost and reproducible behavior because the script predetermines every retrieval step, test execution, and human intervention. Results from ChatRepair, ThinkRepair, Agentless, and KGCompass indicate that repair accuracy increases as the feedback provided to a frozen model becomes more precise; structured traces, knowledge graph context, and reviewer comments offer clearer guidance than raw assertion text. Future work can improve efficiency by prioritizing feedback according to its information value, so that prompts contain only the most helpful facts, and by allowing the loop to increase its analytical depth only when simpler cues are insufficient. Integrating firm control of scripted pipelines with small autonomous routines for complex faults may extend coverage without incurring additional costs. These advances will support the reliable adoption of LLM-based repair in continuous integration practices.